OpenStudy (anonymous):

"without graphing determine the quadrient in which the point (x,y) lies if x<0 and y<0 " A.II B.IV C.I D. III Please help!

OpenStudy (anonymous):

It is on the vertex I think.

OpenStudy (texaschic101):

its the bottom left quadrant...whichever one that is

OpenStudy (anonymous):

It is on three then. texaschic101 is right.

OpenStudy (texaschic101):

I just looked it up...it is 3

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@texaschic101 @Headlessdust41 Thank you! Mind helping me with some other questions?

OpenStudy (anonymous):

Sure.

OpenStudy (texaschic101):

whatcha got

OpenStudy (anonymous):

On which of the following lines does the point (7,1) Lie? A.Y-5x=4 B.Y= X- 10 C.Y= -4x+3 D. Y= -X +8

OpenStudy (anonymous):

Thank you again!

OpenStudy (texaschic101):

we are going to have to sub our points in to see ...(7,1) x = 7 and y = 1 y - 5x = 4 1 - 5(7) = 4 1 - 35 = 4 -34 = 4 (not this one) y = x - 10 1 = 7 - 10 1 = -3 (nope) y = -4x + 3 1 = -4(7) + 3 1 = -28 + 3 1 = - 25 (nope) y = -x + 8 1 = -7 + 8 1 = 1 (It would have to be the last one..lol)

OpenStudy (anonymous):

Last one(: Find the Slope of the line through (-9, -10) and (-2, -5). A.-5/7 B.7/5 C.5/7 D.-7/5

OpenStudy (anonymous):

Thank you so much!!! It means it a lot!

OpenStudy (texaschic101):

slope formula is : slope = (y2 - y1) / (x2 - x1) (-9,-10) x1 = -9 and y1 = -10 (-2,-5) x2 = -2 and y2 = -5 now we sub slope = (-5 - (-10) / (-2 - (-9) slope = (-5 + 10) / (-2 + 9) slope = 5/7 (C) I don't mind helping when I can :)